# CTBC: Contact-Triggered Blind Climbing for Wheeled Bipedal Robots with Instruction Learning and Reinforcement Learning

### v1

@@ Abstract
In recent years, wheeled bipedal robots have gained increasing attention due to their advantages in mobility, such as high-speed locomotion on flat terrain. However, their performance on complex environments (e.g., staircases) remains inferior to that of traditional legged robots. To overcome this limitation, we propose a general contact-triggered blind climbing (CTBC) framework for wheeled bipedal robots. Upon detecting wheel-obstacle contact, the robot triggers a leg-lifting motion to overcome the obstacle. By leveraging a strongly-guided feedforward trajectory, our method enables the robot to rapidly acquire agile leg-lifting skills, significantly enhancing its capability to traverse unstructured terrains. The approach has been experimentally validated and successfully deployed on LimX Dynamics’ wheeled bipedal robot, Tron1. Real-world tests demonstrate that Tron1 can reliably climb obstacles well beyond its wheel radius using only proprioceptive feedback.
@@ II-A RL-Based Legged Locomotion
In 2019, Hwangbo et al. [5] introduced reinforcement learning to legged robot control by proposing a policy conditioned on desired velocity that outputs joint-position targets, augmented with an actuator network to enhance motor modeling accuracy. Building on this framework, Lee et al. [6] developed a teacher–student architecture that enables robust traversal of hills, steps, and other challenging terrains using only proprioception. Siekmann et al. [7] further demonstrated that an RL policy trained solely on proprioception can drive the bipedal robot Cassie to blindly ascend stairs. Leveraging GPU-parallel simulation in Isaac Gym, Rudin et al. [8] trained a locomotion policy with Proximal Policy Optimization (PPO) in just 20 minutes that transfers zero-shot to real hardware. Extending this work, Rudin et al. [9] introduced a task formulation based on positional goals: within a time limit, the robot must reach a target location, autonomously planning both path and motion to overcome obstacles and complete navigation without additional motion priors.

### v2

@@ Abstract
In recent years, wheeled bipedal robots have gained increasing attention due to their advantages in mobility, such as high-speed locomotion on flat terrain. However, their performance on complex environments (e.g., staircases) remains inferior to that of traditional legged robots. To overcome this limitation, we propose a general contact-triggered blind climbing (CTBC) framework for wheeled bipedal robots. Upon detecting wheel-obstacle contact, the robot triggers a leg-lifting motion to overcome the obstacle. By leveraging a strongly-guided feedforward trajectory, our method enables the robot to rapidly acquire agile leg-lifting skills, significantly enhancing its capability to traverse unstructured terrains. The approach has been experimentally validated and successfully deployed on LimX Dynamics’ wheeled bipedal robot, Tron1. Real-world tests demonstrate that Tron1 can reliably climb obstacles well beyond its wheel radius using only proprioceptive feedback. Project page: https://ctbc-for-wheeled-bipedal-robots.github.io/
@@ II-A RL-Based Legged Locomotion
In 2019, Hwangbo et al. [5] introduced reinforcement learning to legged robot control by proposing a policy conditioned on desired velocity that outputs joint-position targets, augmented with an actuator network to enhance motor modeling accuracy. Building directly on this seminal framework, Lee et al. [6] developed a teacher–student architecture that enables robust traversal of hills, steps, and other challenging terrains using only on-board proprioception. Siekmann et al. [7] further demonstrated that an RL policy trained solely on proprioception can drive the bipedal robot Cassie to blindly ascend real-world stairs. Leveraging massively GPU-parallel simulation in Isaac Gym, Rudin et al. [8] trained a locomotion policy with Proximal Policy Optimization (PPO) in just 20 minutes that transfers zero-shot to real hardware. Extending this promising work, Rudin et al. [9] introduced a task formulation based on positional goals: within a strict time limit, the robot must reach a target location, autonomously planning both path and motion to overcome obstacles and complete navigation without additional motion priors.